OpenStudy (anonymous):

What is the circumference of a circle with an area of 64pi ? a) 8pi b) 32pi c) 64pi d) 16pi

OpenStudy (anonymous):

Area = Pi*r^2 = 64*pi Cancle the pi on both sides, you have: r^2 = 64 Take square root, r = 8. Circumference = 2*r*pi Circumference = 2*8*pi = ?